# Bayesian network 3D event reconstruction in the Cygno optical TPC for dark matter direct detection

### v1

@@ 5 Fitting strategy and dataset
The Bayesian network presented in Eq. 5 and Fig. 5 can be implemented in three different configurations, depending on the set of parameters to be inferred.
5.1 Calibration of the PMT response
A first application of the model is the calibration of the PMT response, i.e., the determination of the calibration coefficient $C_{i}$ for each PMT. This can be achieved by analyzing events originating from a known position in the $(XY)$ plane, with identical energy deposition $L$ and track topology. To this end, a dataset is selected consisting of events with a single light cluster in the camera image and a single signal recorded by each of the four PMTs. These events are acquired during exposure to a 55Fe radioactive source, which emits both $K_{\alpha}$ and $K_{\beta}$ X-rays of 55Mn, though for our detector these lines merge into an effectively monochromatic line averaged at 5.9 keV due to its limited resolution. A clean sample of 55Fe events is selected following the analysis reported in previous studies LIME ; bib:fe55 .
The PMT waveforms produced by the resulting electron recoils typically exhibit a single peak, as shown in Fig. 3. The integrated charge is computed within a 60 ns time window centered on the peak. For the Bayesian fit, a flat prior is assigned to each calibration parameter, defined over a positive range bounded from above at a value of 107 a.u. just to preserve the unitary of the pdf.
@@ 5.2 Reconstruction of localized tracks
Once the calibration constants are fixed (e.g., using the values obtained in the previous step), the Bayesian framework can be employed to reconstruct the position and intensity of localized tracks. For each set of four PMT signals, the parameters are inferred using Eq. 5, assuming $N=1$ source. Flat priors are assigned to $X$ and $Y$, constrained within the GEM plane (33 $\times$ 33 cm2), while the prior on light intensity is flat over positive values below a predefined upper bound. For this task we use events collected during exposure to 55Fe radioactive source.
@@
Figure 9(a) shows the spatial distribution of the spots reconstructed with the camera-based algorithm and figure 9(b) the PMT-based algorithm. The distribution appears narrower along the x-axis due to the presence of a collimator in front of the source. As a result, the spot density is higher along X, which – as will be discussed later – leads to better spatial resolution in that direction compared to Y. In addition, figure 9(c) shows that the reconstructed energies from the two sub-detectors are in very good agreement, demonstrating that the limiting factor for the energy resolution is the number of photons produced at the end of the charge amplification process, rather than the readout method.
Figure 9: Planar distribution of the tracks in the camera field of view reconstructed using (9(a)) the APS-sCMOS analysis, and (9(b)) the PMT-based Bayesian algorithm, using a dataset acquired in with a 55Fe radioactive source positioned above the detector. (9(c)) Reconstructed energy spectrum with both analysis.
Events featuring a single visible localized track and a single PMT trigger are used to evaluate the accuracy of the reconstruction. Fig. 10 shows the distribution of the residuals $\Delta X,\,\Delta Y$ expressed as mean and standard deviation between the PMT-based and camera-based reconstructions of the $X$ and $Y$ coordinates, respectively:
$\Delta X$ $=(-0.07\pm 0.85)~{}\text{cm}$ (7)
$\Delta Y$ $=(-0.2\pm 1.6)~{}\text{cm}$ (8)
The uncertainty of these values represent the $(X,Y)$ spatial resolution achievable by the PMT-only approach.
In addition, a toy Monte Carlo simulation was developed to validate the performance of the reconstruction algorithm. By generating synthetic spot-like emissions with known positions and intensities, and applying the Bayesian reconstruction framework, we confirmed that the algorithm accurately recovers both the source positions and the emitted light within the expected uncertainties, further supporting the reliability of the PMT-based approach.
@@ 5.3 Extended tracks
The BAT fit infers the $(X,Y,L)$ parameters of the ionization tracks. This information is combined with the $\Delta Z$ component extracted from the analysis of PMT signals, which – under the assumption of strictly straight tracks, such as MIP-like particles, alpha particles, and nuclear recoils – corresponds to the time extension of the PMT signal. This approach enables a full 3D reconstruction of the particle’s trajectory. The $(X,Y)$ coordinates inferred from the BAT fit are then used to associate each PMT signal with a corresponding track in the camera image, where the transverse spatial resolution is significantly higher.
@@
For events producing extended ionization trails, the reconstruction procedure using the BAT fit follows the same approach as that adopted for localized interactions (see Section 5.2). When the PMT signal exhibits multiple peaks – typical of MIP-like particles – a peak-finding algorithm is applied to identify the dominant features. Only peaks that are observed within the same time window by at least two PMTs, exceeding a defined threshold, and separated by a minimum of 60 ns are retained. Each selected peak is then treated as an independent localized event and reconstructed using the method described in Section 5.2. The $\Delta Z$ component of each reconstructed segment is obtained by measuring the time difference between consecutive peaks and converting it into a spatial distance using the electron drift velocity.
@@
In the case of alpha particles, where the PMT signal appears as a continuous step-like signal without prominent peaks, a different strategy is adopted. The signal is divided into short time windows, each approximately 60 ns wide, corresponding to the typical duration of a localized interaction. The integrated charge in each time slice is computed and fitted using the BAT algorithm, resulting in a set of $(X,Y,L)$ points. This information can be used either to match the PMT signals with the corresponding pixel cluster in the camera image, retrieving the $(XY)$ projection, or to directly reconstruct the 3D shape of the alpha track. Since individual peaks are absent in this case, the $\Delta Z$ coordinate is extracted by measuring the full time-over-threshold width of the signal.
@@
An example of this procedure is shown in figure 11, illustrating the key steps. In panel 11(a), the alpha PMT signals are displayed, showing a high-amplitude, sustained signal with no prominent peaks. The multiple time windows used to segment the signal into short, localized interactions (60 ns each) are also indicated. The resulting posterior $(X,Y)$ positions inferred with the BAT fit are shown as red stars in panel 11(b), overlaid on the corresponding camera image. A clear match between the PMT-based and camera-based reconstructions is observed. The time-over-threshold for this track is measured to be 400 ns, which corresponds to $\Delta Z$ = 2.2 cm, assuming the nominal drift velocity. By combining the information from the PMT signals and the camera image, a full 3D reconstruction of the alpha track is obtained, as shown in panel 11(c). For improved visualization, an ionization cloud is rendered by sampling random points from the transverse light distribution.
@@
Figure 11: 3D reconstruction of an extended alpha particle track. (11(a)) PMT signals with highlighted time windows used for the BAT fit; (11(b)) overlay of the BAT-reconstructed positions (red stars) on the corresponding camera image; (11(c)) final 3D representation of the alpha track combining PMT and camera information.
To assess the accuracy of the matching between the track reconstructed from PMT signals and the one observed in the camera image, the following procedure is applied: the track identified in the image is resampled to match the number of points in the corresponding PMT waveform. A point-by-point distance is then computed between the two tracks, serving as a first-order metric of the matching performance. The residuals in the $\Delta X$ and $\Delta Y$ directions are shown in Figure 12. Compared to the results obtained for spot-like (localized) interactions in Figure 10, the matching residuals exhibit a slightly larger standard deviation, with an average of 124 pixels, corresponding to 1.9 cm. The residuals show an offset along the $X$-direction, likely due to a systematic effect introduced by the method used to compute the distance between the two views of the track (APS-sCMOS and PMT). Further investigation is ongoing to understand this issue. The comparison between the reconstructed PMT track and the image in the $(XY)$ plane shows a good agreement in both direction and topology, typically within a few degrees. Since the camera does not provide $Z$ information, we cannot directly validate the angular accuracy in three dimensions. However, the consistency observed in $(X,Y)$ suggests that the resolution in $Z$ is of comparable quality, given the similar level of PMT signal constraint and the use of a common light emission model. This supports the robustness of the 3D directional reconstruction based solely on PMT data. The current model does not resolve the track sense (head-tail), as it assumes a symmetric light emission along the track. This could be improved in future work by including asymmetric light yield models.
@@
The results presented for extended tracks are promising in the framework of the 3D reconstruction of particles in CYGNO detectors. Additional studies are being carried out concerning the optimization of these results and application to other types of particles. In an upcoming paper, we further explore this technique, and improve upon it, by retrieving not only the 3D shape of the track, but also its direction and sense, a fundamental information for directional dark matter searches. This is used to study the LIME alpha background, and in particular the presence of Radon progeny recoils.
@@ 6 Conclusions
This work demonstrates, for the first time, the feasibility of full 3D track reconstruction using only PMT signals in a gaseous TPC with optical readout. By modeling the light collection process probabilistically, our method enables robust spatial localization and energy estimation directly from time-resolved scintillation data. This approach is validated using data from the CYGNO LIME prototype and successfully reconstructs both localized and extended ionization tracks with sub-centimeter precision.
When combined with the high-resolution 2D imaging provided by the APS-sCMOS camera, the Bayesian reconstruction framework further improves spatial and energy resolution, allowing for precise 3D mapping of ionizing events. The method is implemented using the Bayesian Analysis Toolkit (BAT) and Metropolis-Hastings MCMC sampling, and has proven effective both for PMT calibration and event-by-event reconstruction.
These results represent a significant advancement for optical TPCs in the context of rare-event searches. In particular, the ability to reconstruct 3D trajectories without relying on a pixelated readout opens the door to scalable and cost-effective designs for directional dark matter detection.
While the current implementation assumes a single extended light source and does not account for optical effects like refraction through different materials along the path, it already proves effective across a variety of track topologies. Future work will address scenarios with low light yield or overlapping signals, and will further refine the model to enhance robustness and extend its applicability to rare-event searches.
Ongoing developments aim to extract not only the track geometry but also its direction and sense, a key requirement for identifying WIMP-induced nuclear recoils. These techniques are currently being applied to the study of alpha-induced backgrounds in LIME, including Radon progeny recoils, and will be detailed in a forthcoming publication.
[…]
In conclusion, the methodology developed in this work lays the foundation for a new generation of optically read out TPCs, offering high-resolution, 3D directional sensitivity – an essential capability in the future of dark matter detection.
Acknowledgement

### v2

@@ 5 Fitting strategy and dataset
The Bayesian network of Eq. 8 and Fig. 6 can be implemented in three configurations: (1) PMT calibration, where the fit runs on $n$ point-like events of known light $L_{j}$ and position $(X_{j},Y_{j})$ (thus $N=n$ in Eq. 8) to determine the calibration constants $C_{i}$ (Sec. 5.1); (2) reconstruction of spot-like events to infer $L_{j}$ and $(X_{j},Y_{j})$ (Sec. 5.2); (3) reconstruction of extended tracks to infer $L_{j}$ and $(X_{j},Y_{j})$ of their sub-segments (Sec. 5.3). For modes (2) and (3), the fit is performed with $N=1$ in Eq. 8.
5.1 In-situ calibration of the PMT response
A first application of the model is the in-situ calibration of the PMT response, whose relevance lies in accounting not only for gain equalization but also for geometric and alignment effects, as well as material-dependent factors influencing light collection. The calibration consists in the determination of the coefficient $C_{i}$ for each PMT.
This is achieved by analyzing events from a known position in the $(XY)$ plane with identical energy deposition $L$ and track topology. For this purpose, we select a dataset of $N=n$ events characterized by a single light cluster in the camera image and a single signal recorded by each of the four PMTs. These events are acquired during exposure to a 55Fe radioactive source similarly to what described in section 2.2.
@@ 5.2 Reconstruction of localized tracks
Once the calibration constants are fixed (e.g., using the values obtained in the previous step), the Bayesian framework can be employed to reconstruct the position and intensity of localized tracks. For each set of four PMT signals, the parameters are inferred using Eq. 8, with $N=1$ source. Flat priors are assigned to $X$ and $Y$, constrained within the GEM plane (33 $\times$ 33 cm2), while the prior on light intensity is flat over positive values below a predefined upper bound. For this task we use events collected during exposure to 55Fe radioactive source.
@@
Figure 10(a) shows the spatial distribution of the spots reconstructed with the camera-based algorithm and figure 10(b) the PMT-based algorithm. The distribution appears narrower along the x-axis due to the presence of a collimator in front of the source. As a result, the spot density is higher along X, which – as will be discussed later – leads to better spatial resolution in that direction compared to Y. In addition, figure 10(c) shows that the reconstructed energies from the two sub-detectors are in very good agreement. For the CMOS, the energy estimation is obtained from the total photon counts in the pixels constituting the tracks, while for the PMT it comes directly from the BAT-fit light output variable $(L)$. A single-point calibration is applied by associating the main peak in both sensors distributions to the 55Fe peak of 5.9 keV. The observed non-gaussian tails are due to poorly reconstructed tracks and background events. This result demonstrates that the limiting factor for the energy resolution is the number of photons produced at the end of the charge amplification process, rather than the readout method.
Figure 10: Planar distribution of the tracks in the camera field of view reconstructed using (10(a)) the APS-sCMOS analysis, and (10(b)) the PMT-based Bayesian algorithm, using a dataset acquired in with a 55Fe radioactive source positioned above the detector. (10(c)) Reconstructed energy spectrum with both analysis.
Events featuring a single visible localized track and a single PMT trigger are used to evaluate the reconstruction accuracy. Figure 11 shows the distributions of the residuals $\Delta X$ and $\Delta Y$ between the PMT-based and camera-based reconstructions of the $X$ and $Y$ coordinates, respectively. The mean values and standard deviations are
$\Delta X$ $=(-0.05\pm 0.81)~\text{cm},$ (9)
$\Delta Y$ $=(-0.1\pm 1.5)~\text{cm}.$ (10)
The quoted uncertainties represent the $(X,Y)$ spatial resolution achievable with the PMT-only approach. The difference between the two coordinates reflects the $X$-axis collimation of the source, which results in a different spatial spread of the events used for this study (see Figs. 10(a) and 10(b)).
In addition, a toy Monte Carlo simulation was developed to validate the internal consistency of the reconstruction algorithm. By generating synthetic spot-like emissions with known positions and intensities, and applying the Bayesian reconstruction framework, we confirmed that the algorithm accurately recovers both the source positions and the emitted light within the expected uncertainties, further supporting the reliability of the PMT-based approach.
@@ 5.3 Extended tracks
We emphasize that even in the case of extended tracks, the reconstruction is carried out under the assumption of local straightness, in line with the current scope of the work. This choice ensures consistency with the nuclear-recoil topology of interest, while laying the groundwork for future extensions to more general track shapes.
@@
The BAT fit infers the $(X,Y,L)$ parameters of the ionization tracks. This information is combined with the $\Delta Z$ component extracted from the analysis of PMT signals which, under the assumption of strictly straight tracks – such as MIP-like particles, alpha particles, and nuclear recoils – enables a 3D reconstruction. The $(X,Y)$ coordinates inferred from the BAT fit are then used to associate each PMT signal with a corresponding track in the camera image, where the transverse spatial resolution is significantly higher.
Events producing extended ionization trails are first reduced to point-like tracklets through a slicing procedure. Each individual tracklet is then reconstructed using the BAT fit, as is done for localized interactions (see Section 5.2). Consequently, the computational cost for extended tracks is the same as that for localized ones, scaled by the number of slices.
@@
When the PMT signal exhibits multiple peaks – typical of MIP-like particles and extended ER – a peak-finding algorithm is applied to identify the dominant features. Only peaks that are observed within the same time window by at least two PMTs, exceeding a defined threshold and separated by a minimum of 60 ns, are retained. Each selected peak is then treated as an independent localized event and reconstructed using the method described in Section 5.2. The $\Delta Z$ component of each reconstructed segment is obtained by measuring the time difference between consecutive peaks and converting it into a spatial distance using the electron drift velocity.
@@
In the case of alpha particles, where the PMT signal appears as a continuous step-like signal without prominent peaks, a different strategy is adopted. The signal is divided into short time windows, each 60 ns wide, corresponding to the typical duration of a localized interaction. The integrated charge in each time slice is computed and fitted using the BAT algorithm, resulting in a set of $(X,Y,L)$ points. This information can be used either to match the PMT signals with the corresponding pixel cluster in the camera image, retrieving the $(XY)$ projection, or to directly reconstruct the 3D shape of the alpha track. Since individual peaks are absent in this case, the $\Delta Z$ coordinate is extracted by measuring the time difference between the last and the first slice. An example of this procedure is shown in Fig. 12, illustrating the key steps. In panel 12(a), the alpha PMT signals are displayed, showing a high-amplitude, sustained signal with no prominent peaks.
@@
The multiple time windows used to segment the signal into short, localized interactions (60 ns each) are also indicated. This segmentation assumes straight tracks, under which condition a finer subdivision in $(x,y)$ would ultimately need to be recombined before matching with the PMT slices, yielding the same effective result. The resulting posterior $(X,Y)$ positions inferred with the BAT fit are shown as red stars in panel 12(b), overlaid on the corresponding camera image. A clear match between the PMT-based and camera-based reconstructions is observed. The time-length for this track is measured to be 400 ns, which corresponds to $\Delta Z$ = 2.2 cm, assuming the nominal drift velocity. By combining the information from the PMT signals and the camera image, a 3D reconstruction of the alpha track is obtained, as shown in panel 12(c), illustrating how the information from the two sensors can be merged into a single event display, which is ultimately also one of the objectives of CYGNO. For improved visualization, an ionization cloud is rendered by sampling random points from the transverse light distribution.
@@
Figure 12: 3D reconstruction of an extended alpha particle track. (12(a)) PMT signals with highlighted time windows used for the BAT fit; (12(b)) overlay of the BAT-reconstructed positions (red stars) on the corresponding camera image; (12(c)) final 3D representation of the alpha track combining PMT and camera information.
@@
To quantify the matching between the PMT-based reconstruction and the track observed in the camera, the image track is resampled to the same number of points as the PMT waveform. Edge pixels are identified as start and end points, and intermediate points are uniformly distributed along the track in the $(XY)$ plane. A point-by-point distance is then computed between the two reconstructions. The residuals in $\Delta X$ and $\Delta Y$ are reported in Fig. 13. A systematic offset in the $X$ direction is observed, likely due to the distance-metric definition; further investigation is ongoing. Compared to the localized interactions of Fig. 11, the residuals show a non Gaussian shape and a slightly larger spread of 1.9 cm. This is justified as the fit performance depends on the $(X,Y)$ position of the track and multiple correlated points per track contribute to the histograms. Event-by-event systematics, combined with the non-isotropic alpha emission in the detector, naturally lead to the observed non-Gaussian features and asymmetries. Overall, the reconstructed PMT tracks match the image projection in direction and topology. Since the camera provides no $Z$ information, the angular accuracy in three dimensions cannot be directly validated, but the consistency in $(XY)$ suggests comparable resolution along $Z$. This supports the robustness of the 3D reconstruction from PMT data alone.
The current model does not resolve the track sense (head–tail), as it assumes symmetric light emission. This limitation could be overcome by incorporating asymmetric light-yield models in future work. In a forthcoming paper we will extend the method to retrieve not only the 3D shape, but also the direction and sense of tracks – crucial information for directional dark matter searches. This will also be applied to the study of alpha backgrounds in LIME, in particular radon progeny recoils.
@@ 6 Conclusions
To our knowledge, this work provides the first demonstration that PMT signals alone can be used to reconstruct both the three-dimensional topology and the energy of ionization tracks in a gaseous TPC with optical readout. By modeling light collection probabilistically within the BAT framework and using MCMC sampling, we achieve robust in-situ calibration and event-by-event reconstruction. Validation with data from the CYGNO LIME prototype shows that both localized and extended tracks can be reconstructed with sub-centimeter precision.
This represents a significant step forward for optical TPCs in rare-event searches. The ability to perform 3D reconstruction without a pixelated readout paves the way for scalable, cost-effective designs for directional dark matter detection. While the present implementation assumes straight tracks and neglects optical effects such as refraction at material boundaries, it already proves effective across diverse topologies. Future improvements will target low-light and overlapping signals, as well as asymmetric light-yield models to recover track direction and sense, crucial for identifying WIMP-induced nuclear recoils.
These techniques are currently being applied to the study of alpha-induced backgrounds in LIME, including radon progeny recoils, and will be detailed in forthcoming work. In parallel, the results reported here have informed the optimization of PMT layout for CYGNO-04, a $\mathcal{O}$(1 m3) detector under design.
In conclusion, the PMT-only reconstruction with BAT achieves a sub-centimeter spatial resolutions and an energy resolutions comparable to that of the sCMOS, enables robust event matching even under pile-up, and provides a solid basis for 3D directional sensitivity, an essential capability for future dark matter searches.
Acknowledgements